# Large Language Models Lack Understanding of Character Composition of Words

### v1

@@ 1 Introduction
Humans are able to instantly recognize which characters constitute a given word. However, large language models, most of which are trained at token-level, struggle to grasp the nuances of character composition within words. This difficulty arises from the fact that LLMs primarily learn at the token level, where words are treated as indivisible units separated by spaces or punctuation marks. Consequently, LLMs lack the fine-grained understanding of character-level relationships and morphology that humans possess. Understanding character composition is crucial for various linguistic tasks, including morphological analysis, semantic interpretation, and language generation. As such, addressing the challenge of character composition is essential for enhancing the reliability and performance of LLMs across a diverse range of languages and writing systems.
@@ 2 Related Works
Subsequent works in language modeling have further explored the integration of character-level information. For instance, (Peters et al., 2018) introduced deep contextualized word representations (ELMo), which enhance word embeddings by considering the internal structure of words through character-level convolutions. This method significantly improved the performance of various NLP tasks by capturing complex word morphologies. Additionally, (Akbik et al., 2019) proposed Flair embeddings, which combine character-level embeddings with contextual string embeddings to provide a more comprehensive representation of words in their context. Despite these advancements, there remains a gap in specifically addressing character composition understanding. (Clark et al., 2020) introduced ELECTRA, a pre-training method that includes a discriminative component to identify corruptions at the token level, which indirectly benefits from finer-grained text representations. However, the primary focus has been on token-level tasks rather than explicit character composition understanding.
@@ 3.1 Setting
We experimented with 4 publicly available LLMs, namely GPT4 (Achiam et al., 2023), Claude (Claude, 2023), Gemini 1.5 (Reid et al., 2024), and Mistral 7B (Jiang et al., 2023). We randomly sampled words, phrases, or sentences from Wikipedia corpus. Note that, while it is possible that such publicly available text was used during the pre-training of target LLMs, the character-based nature of our experiments prevents the models from taking advantage of it, and indeed, the results in Sec 3.2 seem to reinforce the claim. For each task, 100 prompts were used, where each prompt may contain multiple answers. In order to compare the LLM’s understanding of character composition with that of humans, we also asked human annotators to perform exactly the same tasks, providing identical prompts and passages.
In order to compare LLMs’ performances at character level and token level tasks, we also extend each task described above to token level tasks. Word retrieval is extended to sentence retrieval, where the model is given 5-sentence passage and is asked to return all sentences containing a target word. Insertion and deletion work similarly by providing target word and position within sentence, whereas we provide target word and another input word for replacement task. Reordering and counting are extended similarly. For reordering, as with character-level reordering, we only compute accuracy from whether the final answer is correct, without computing precision and recall for each reordered word.
@@ 3.2 Results
Table 1 summarizes the results of our experiments with precision, recall, and F-score for each task at character level. For token level, we only report F-score for brevity in Table 2. It is clearly shown that, for most tasks, all target LLMs display severely degraded performance at character level when compared to token level. While discrepancies exist among respective models’ performances, none of them rises to the level of demonstrating a clear superiority over other models. It is also out of scope of this paper to determine which LLM is better, as our focus is on assessing LLMs in terms of understanding character composition in general.
Humans, not surprisingly, demonstrated near-perfect performance throughout all tasks. There was hardly any mistake in precision, while defects in recall tended to occur mostly around characters that are placed in the middle of the word, rather than beginning or the end, suggesting attention to saliency in human perception of character composition. Considering that humans have been surpassed by LLMs in many NLP tasks that are supposedly much more complex, our results suggest an unsettling dichotomy between LLM’s capability at token-level and character-level tasks.
Table 3 shows some of the failure cases for each model at character level. It is notable that the tasks for which LLMs struggled the most frequently involved specifying positions of the characters, mostly using numbers, as in insertion or deletion tasks. It should be noted that a similar performance decline was observed even at token level, as illustrated in Table 2. Table 4 shows example failure cases at token level. This suggests that some of the limitations in understanding character composition may not simply be attributed to the fact that LLMs are trained at token level, but to a more fundamental drawback in their training approach in general.
@@
Notably, all LLMs performed far better on character reordering task than on other tasks, closely trailing the performance of humans. We conjecture that this is due to abundant resources available online about anagram, which are likely to have been used in pre-training of the models. Even when the newly formed words are non-existing words, many of them are likely to have appeared in the training corpora as possible anagrams of an existing word . It is therefore only natural that all models struggled with character reordering as the word gets longer, or with an unknown word, as shown in Table 4.
@@ 4 Discussion
A promising direction to address this limitation involves embedding character-level information directly into word embeddings, enabling models to capture the intricate relationships and structures within individual characters. For example, BERT (Devlin et al., 2019) represents input tokens not only with token embedding, but also with segment embedding, which indicates the sentence that the token belongs to, and position embedding, which shows the position of the token within the sentence. A similar structural approach can be made with respect to character, where character is embedded also with information of the word it belongs to, and its position within the word. Such multi-level embedding strategy could significantly enhance the model’s ability to understand and manipulate text at a finer granularity. Furthermore, leveraging achievements in subword tokenization methods, such as Byte Pair Encoding (BPE) (Sennrich et al., 2015), which breaks down words into subword units, can complement the multi-level embedding approach. Such dual-layered approach can help ensure that the model obtains a robust understanding of word composition while being sensitive to the arrangement of characters within words.
Another potential line of approach involves harnessing visual recognition techniques to simulate human-like character perception. In scene text recognition literature, there has been a number of endeavors to integrate computer vision methodologies to visually identify characters, replicating the cognitive processes humans employ when reading and comprehending text (Du et al., 2022; Bartz et al., 2017). By leveraging the complementary strengths of both domains, these approaches may potentially offer novel opportunities for improving robustness for character-level comprehension within large language models.

### v2

@@ 1 Introduction
Humans are able to instantly recognize which characters constitute a given word. However, large language models, most of which are trained at token-level, struggle to grasp the nuances of character composition within words. This difficulty arises from the fact that LLMs primarily learn at the token level, where words are treated as indivisible units separated by spaces or punctuation marks. Consequently, LLMs lack the fine-grained understanding of character-level relationships and morphology that humans possess. Understanding character composition is crucial for various linguistic tasks, including morphological analysis, semantic interpretation, and language generation. As such, addressing the challenge of character composition is essential for enhancing the reliability of LLMs across a diverse range of languages and writing systems.
@@ 2 Related Works
Subsequent works in language modeling have further explored the integration of character-level information. For instance, (Peters et al., 2018) introduced deep contextualized word representations (ELMo), which enhance word embeddings by considering the internal structure of words through character-level convolutions. This method significantly improved the performance of various NLP tasks by capturing complex word morphologies. (Akbik et al., 2019) proposed Flair embeddings, which combine character-level embeddings with contextual string embeddings to provide a more comprehensive representation of words in their context. (Clark et al., 2020) introduced ELECTRA, a pre-training method that includes a discriminative component to identify corruptions at the token level, which indirectly benefits from finer-grained text representations. For most of these works, however, the primary focus has been on token-level tasks rather than specifically addressing character composition understanding.
With regards to more recent LLMs, there have been a number of works that highlight their downsides from various angles. For example, (Qian et al., 2022) claims that LLMs struggle with arithmetic and symbolic manipulations, while (Lee & Lim, 2024) shows that LLMs fail to learn physical manifestation of language, such as the visuals and sounds of the language. (Truong et al., 2023) also shows that LLMs’ performances degrade when negation is involved. With regards to the character composition, there have been a few attempts to benchmark the performances of LLMs (Srivastava et al., 2022; Efrat et al., 2022), although the scope of evaluating character composition was highly restricted, with stronger emphasis on evaluation of word-level understanding.
@@ 3.1 Setting
We experimented with 4 publicly available LLMs, namely GPT4 (Achiam et al., 2023), Claude (Claude, 2023), Gemini 1.5 (Reid et al., 2024), and Mistral 7B (Jiang et al., 2023). We randomly sampled words, phrases, or sentences from Wikipedia corpus. Note that, while it is possible that such publicly available text was used during the pre-training of target LLMs, the character-based nature of our experiments prevents the models from taking advantage of it, and the results in Sec 3.2 seem to reinforce the claim. For each task, 100 prompts were used, where each prompt may contain multiple answers. In order to compare the LLM’s understanding of character composition with that of humans, we also asked human annotators to perform exactly the same tasks, providing identical prompts and passages.
In order to compare LLMs’ performances at character level and token level tasks, we also extend each task described above to token level tasks. Word retrieval is extended to sentence retrieval, where the model is given 5-sentence passage and is asked to return all sentences containing a target word. Insertion and deletion work similarly by providing target word and position within sentence, whereas we provide target word and another input word for replacement task. Reordering and counting are extended similarly. For reordering, as with character-level reordering, we compute accuracy from whether the final answer is correct, without computing precision and recall for each reordered word.
@@ 3.2 Results
Table 1 summarizes the results of our experiments with precision, recall, and F-score for each task at character level. For token level, we only report F-score for brevity in Table 2. It is clearly shown that, for most tasks, all target LLMs display severely degraded performance at character level when compared to token level. While discrepancies exist among respective models’ performances, none rises to the level of demonstrating a clear superiority over other models. It is also out of scope of this paper to determine which LLM is better, as our focus is on assessing LLMs in terms of understanding character composition in general.
Humans, not surprisingly, demonstrated near-perfect performance throughout all tasks. There was hardly any mistake in precision, while defects in recall tended to occur mostly around characters that are placed in the middle of the word, rather than beginning or the end, suggesting attention to saliency in human perception of character composition. Considering that humans have been surpassed by LLMs in many NLP tasks that are supposedly more complex, our results suggest an unsettling dichotomy between LLM’s capability at token-level and character-level tasks.
Table 3 shows some of the failure cases for each model at character level. It is notable that the tasks for which LLMs struggled the most frequently involved specifying positions of the characters, mostly using numbers, as in insertion or deletion tasks. It should be noted that a similar performance decline was observed even at token level, as illustrated in Table 2. Table 4 shows example failure cases at token level. This suggests that some of the limitations in understanding character composition may not simply be attributed to the fact that LLMs are trained at token level, but to a more fundamental drawback in their training approach. Notably, all LLMs performed far better on character reordering task than on other tasks, closely trailing the performance of humans. We conjecture that this is due to abundant resources available online about anagram, which are likely to have been used in pre-training of the models. Even when the newly formed words are non-existing words, many of them are likely to have appeared in the training corpora as possible anagrams of an existing word . It is therefore only natural that all models struggled with character reordering as the word gets longer, or with an unknown word, as shown in Table 4.
@@ 4 Discussion
A promising direction to address this limitation involves embedding character-level information directly into word embeddings, enabling models to capture the intricate relationships and structures within individual characters. For example, BERT (Devlin et al., 2019) represents input tokens not only with token embedding, but also with segment embedding, which indicates the sentence that the token belongs to, and position embedding, which shows the position of the token within the sentence. A similar structural approach can be made with respect to character, where character is embedded also with information of the word it belongs to, and its position within the word. Such multi-level embedding strategy could significantly enhance the model’s ability to understand and manipulate text at a finer granularity, and can help ensure that the model obtains a robust understanding of word composition while being sensitive to the arrangement of characters within words. Another potential line of approach involves harnessing visual recognition techniques to simulate human-like character perception. In scene text recognition literature, there has been a number of endeavors to integrate computer vision methodologies to visually identify characters, replicating the cognitive processes humans employ when reading and comprehending text (Du et al., 2022; Bartz et al., 2017). By leveraging the complementary strengths of both domains, these approaches may potentially offer novel opportunities for improving robustness for character-level comprehension within large language models.